Will someone please help me! I do not understand how to get to the polar form!
Write the following as a complex number in polar form:
sqrt.3 + i
Thank you for your time, Natalie
: Will someone please help me! I do not understand how to get to the polar form!
Write the following as a complex number in polar form:
sqrt.3 + i
Thank you for your time, Natalie

Answer by scott8148(2724)

with complex numbers, the real part is on the horizontal (x) axis
__ and the imaginary part is on the vertical (y) axis

so sqrt(3) on the x and 1 on the y forms two sides of a right triangle
__ the 3rd side is the hypotenuse; which is the polar radius

using Pythagoras, r^2=x^2+y^2 __ r^2=[sqrt(3)]^2+1^2 __ r^2=3+1 __ r=2

the tangent of the polar angle is [imaginary]/[real] or y/x
__ tan(Θ)=1/sqrt(3) __ Θ=30º or π/6 radians